What is expected annual return on savings before retirement (convert to APR)? (Use nominal function & show formula). Use information below:

How much money do you save each period: $750

Years until retirement: 30 years

Periods in a year: 4

Expected return on your savings before retirement (EAR): 8.00%

Expected return on savings during retirement (APR): 6.00%
Answer #1

Expected return on savings before retirement (EAR): 8.00%

As 1+EAR = (1+APR/no. of compounding periods in a year)^no of compounding periods in a year

=> 1+0.08 = (1+APR/4)^4

=> 1+APR/4 = 1.08^(1/4)

=> APR/4 = 0.0194265

APR =0.0777062

or 7.77%